# Graph Reconstruction with a Connected Components Oracle

### v1

@@ Abstract
    For a hidden graph with $n$ vertices, $m$ edges, maximum degree $\Delta$, and treewidth $k$, GR can be solved in $\mathcal{O}(\min\{m,\Delta^{2},k^{2}\}\cdot\log n)$ CC queries by an adaptive randomized algorithm.
@@ 1 Introduction
The query complexity of GR has been studied for varying oracles and families of graphs. One main line of research has focused on the independent set (IS) oracle that only reports whether the queried subset is an independent set. This oracle has been applied to learning matchings [4, 8], cycles [11], stars [3], cliques [3], and general graphs [1, 6]. Recently, Konrad et al. [16] proposed a more powerful oracle that returns a maximal independent set, and Michel and Scott [19] further refined their lower bounds. Other studied oracles include betweenness [2, 23], distance [7, 13, 18, 22], edge counting [9, 12, 21], and shortest path oracles [21, 24].
@@
In this work, we consider an oracle that returns the set of connected components (CC) on the subgraph induced by the queried set. The CC oracle is thus actually a stronger version of the separation oracle used in learning Markov networks [^1] To answer a $Sep(a,b,S)$ query, one just conducts a $\operatorname{CC}(V\setminus S)$ query and check whether $a$ and $b$ are in the same connected component. Essentially, this means that a $\operatorname{CC}(V\setminus S)$ query answers to $Sep(a,b,S)$ queries for all pairs $a,b\in V\setminus S$.. We therefore study how powerful this CC oracle is as a function of classical graph parameters such as the maximum degree or the treewidth of the hidden graph, when the complexity is measured in terms of the number of performed queries.
@@
Our contributions. We start by devising several algorithms for the GR problem with a CC oracle. A summary of the upper bounds is shown in Table 1. We present results parameterised by three graph parameters: number of edges $m$, maximum degree $\Delta$ and treewidth $k$. With all these parameters, we provide adaptive algorithms (both deterministic and randomized) whose query complexity is polynomial with respect to the parameter and logarithmic with respect to the number of vertices. When the maximum degree of the hidden graph is known, we provide also a non-adaptive algorithm. We also show that algorithms for these three parameters can be combined giving a randomized adaptive algorithm that uses at most $\mathcal{O}(\min\{m,\Delta^{2},k^{2}\}\cdot\log n)$ queries (Theorem 3.27).
We also study lower bounds; a summary is shown in Table 2. We show that it is not possible to construct efficient non-adaptive algorithms. Non-adaptive algorithms require $\Omega(n^{2})$ CC queries in worst-case if no additional information is given (Theorem 4.3). This holds even if we know that the graph has treewidth $2$ and thereby has at most $\mathcal{O}(n)$ edges. There is a peculiar observation that very little adaptivity can help. Algorithm 4 has two rounds of adaptivity and query complexity $\mathcal{O}(k^{2}\log n)$ implying that adding one round of adaptivity reduces the dependence on $n$ from $n^{2}$ to $\log n$.
Furthermore, we show that even adaptive algorithms require $\Omega(m)$, $\Omega(\Delta^{2})$, or $\Omega(k^{2})$ CC queries in the worst-case (Theorem 4.1). Thus, our upper bounds are within a $\log n$ factor from the lower bound.
Finally, we compare the CC oracle with the maximal independent set (MIS) oracle introduced by Konrad et al. [16]. While there are instances where the MIS oracle needs $\mathcal{O}(n)$ queries but $\mathcal{O}(\log n)$ queries are sufficient for the CC oracle (Theorem 5.1), these two oracles seem to be roughly equally powerful in the sense that both upper and lower bounds for $m$ and $\Delta$ are the same for both types of oracles. However, CC oracles and MIS oracles are different and neither of them can be efficiently simulated with the other (Theorems 5.3 and 5.5). We also compare the CC oracle and the separation oracle, showing that simulating the former oracle with the latter one requires a quadratic number of queries in the size of the queried set (Theorem 6.3).
Table 1: Summary of our upper bounds.
@@
Table 2: Summary of our lower bounds, where the Complexity column states that there is an instance on which $\Omega(\cdot)$ CC queries are needed. A lower bound for randomized algorithms holds for deterministic algorithms, and a lower bound for adaptive algorithms holds for non-adaptive algorithms.
@@
Theorem 4.3 $\Omega(n^{2})$ yes no even with $k=2$ and $m=\mathcal{O}(n)$
@@ 2 Notation and Preliminaries
The subgraph of $G$ induced by $S\subseteq V$, denoted by $G[S]$, has the vertex set $S$ and the edge set $\big{\{}uv\in E(G)\colon u,v\in S\big{\}}$. A subset of vertices is an independent set if the subgraph induced by it has no edges. An independent set is maximal if none of its proper supersets are independent.
@@ 3.1 Bounded Number of Edges
Angluin and Chen [6] have proven that $\mathcal{O}(m\log n)$ queries suffice for GR when we have access to an edge-detecting oracle that outputs whether the subgraph induced by the query set $S$ contains any edges. In other words, the edge-detecting oracle is just an IS oracle. A single CC query on $S$ can simulate the IS oracle on $S$, for $S$ is an independent set if and only if all vertices are in distinct connected components. Thus, we get the following corollary.
@@
There is an adaptive deterministic algorithm that solves GR with $\mathcal{O}(m\log n)$ CC queries.
@@ 3.2.1 Derandomization
An $(n,\kappa,\kappa^{2})$-splitter of size $k^{\mathcal{O}(1)}\log n$ can be constructed in time $k^{\mathcal{O}(1)}n\log n$.
@@ 3.3 Bounded Treewidth
Algorithm 2 performs the first step: It outputs a supergraph of the hidden graph. Note that, to get a correct result, one does not actually need to know the treewidth of the hidden graph $G$ but an upper bound of it. However, the number of queries conducted increases if the bound is not tight.
@@ 5   Remove edges uvuv for which $v$ and $u$ are in different components from $G^{+}$;
Next, we will prove the correctness and complexity of Algorithm 2.
@@
We take advantage of the following fact about existence of small balanced separators in bounded treewidth graphs:
@@
A coloring of a graph $G$ is a mapping from the vertex set to some number of colors such that neighboring vertices have always different color. Graphs with small treewidth can be colored with few colors:
@@ 3.3.2 Unknown $k$
There exists an adaptive randomized algorithm that solves GR with $\mathcal{O}(\min\{m,\Delta^{2},k^{2}\}\cdot\log n)$ CC queries. The output of the algorithm is correct with high probability.
@@
We use the adaptive randomized algorithms from Corollary 3.1, 3.15, and 3.25. If ran separately, they would stop after performing some number of queries—say, $t_{1}$, $t_{2}$, and $t_{3}$, respectively. Consider instead running the algorithms in parallel, such that in the $i$th iteration we perform the $i$th query of each of the three algorithms. If any of them would stop after performing that query, we output the graph computed by that algorithm. In total, we thus perform $3\cdot\min\{t_{1},t_{2},t_{3}\}=\mathcal{O}(\min\{\Delta^{2},k^{2},m\}\cdot\log n)$ CC queries. Further, since the output of each algorithm is correct with high probability, so is the output of this algorithm.
@@ 4 Lower Bounds
In this section, we show that the query complexity of our randomized algorithms cannot be improved by a factor larger than $\mathcal{O}(\log n)$ with a proof similar to the lower bound proof of Konrad et al. [16]. Further, we show that non-adaptive algorithms require $\Omega(n^{2})$ CC queries if they are not given additional information about the structure.
@@
No algorithm can solve GR with $o(m)$, $o(k^{2})$ or $o(\Delta^{2})$ CC queries with probability greater than $1/2$.
@@
Suppose now that there is a randomized algorithm that outputs the correct graph for each hidden graph $G\in\mathcal{G}_{\eta}$ with probability greater than $1/2$ while performing at most $\eta^{2}-1$ queries. By Yao’s lemma, there is then a deterministic algorithm that performs $\eta^{2}-1$ queries and outputs the correct graph for more than half of the hidden graphs $G\in\mathcal{G}_{\eta}$. However, observe that for any CC query on a set $S\subseteq V$, there are exactly two possible outputs over the graphs $G\in\mathcal{G}_{\eta}$: the vertices in $S\setminus(C_{1}\cup C_{2})$ are always in connected components of size $1$ and the remaining connected components induced by $S$ are either $S\cap C_{1}$ and $S\cap C_{2}$ or just $S$. Consequently, the deterministic algorithm can identify in $\eta^{2}-1$ queries only $2^{\eta^{2}-1}\leq|\mathcal{G}_{\eta}|/2$ hidden graphs, resulting in a contradiction. Thus, $\Omega(\eta^{2})$ queries are needed. We have $k,\Delta\leq 2\eta$ and $m\leq 2\eta^{2}$ for this family of hidden graphs, completing the proof.
Theorem 4.3.
@@
5 Relationship with the Maximal Independent Set Oracle
In this section, we compare the CC oracle against the maximal independent set (MIS) oracle of Konrad et al. [16], which returns a maximal independent set of $G[S]$. Note that the MIS oracle can be adversarial, that is, it may return the “least useful” maximal independent set. We start this section by demonstrating that CC queries are more powerful than MIS queries on some instances.

### v2

@@ Abstract
    For a hidden graph with $n$ vertices, $m$ edges, maximum degree $\Delta$, and treewidth $k$, GR can be solved in $\mathcal{O}(\min\{m/\log m,\Delta^{2},k^{2}\}\cdot\log n)$ CC queries by an adaptive randomized algorithm.
@@
    For a hidden graph with $n$ vertices and degeneracy $d$, GR can be solved in $\mathcal{O}(d^{2}\log^{2}n)$ CC queries by an adaptive randomized algorithm.
  3. 3.
@@ 1 Introduction
The query complexity of GR has been studied for varying oracles and families of graphs. One main line of research has focused on the independent set (IS) oracle that only reports whether the queried subset is an independent set. This oracle has been applied to learning matchings [4, 8], cycles [13], stars [3], cliques [3], and general graphs [1, 6]. Other studied oracles include betweenness [2, 26], distance [7, 15, 21, 25], edge counting [10, 14, 24], shortest path oracles [24, 28], and connected component counting oracle [9].
From an information-theoretical point of view, many of the previous oracles give only few bits of information per query, and thus struggle in reconstructing larger families of graphs in a small number of queries. Konrad et al. [18] therefore recently proposed a more powerful oracle that returns a maximal independent set. Notably, their algorithms are able to reconstruct hidden graphs with bounded maximum degree in a logarithmic number of queries in the number of vertices because of the increased amount of information the oracle provides. The lower bounds for the query complexity with the oracle were later refined by Michel and Scott [22].
In this paper, we study the connected components (CC) oracle, which returns the sets of vertices in each connected component of the subgraph induced by the queried set. This is thus a stronger version of the connected component counting (#\#$C$) oracle of Black et al. [9], which only returns the number of the connected components. One can also see the CC oracle as a stronger version of the separation oracle used in learning Markov networks [^1] To answer a $\operatorname{Sep}(a,b,S)$ query, one just conducts a $\operatorname{CC}(V\setminus S)$ query and check whether $a$ and $b$ are in the same connected component. Essentially, this means that a $\operatorname{CC}(V\setminus S)$ query answers to $\operatorname{Sep}(a,b,S)$ queries for all pairs $a,b\in V\setminus S$.. We then investigate how powerful this CC oracle is as a function of classical graph parameters such as the maximum degree, the treewidth, and the degeneracy of the hidden graph, when the complexity is measured in terms of the number of performed queries.
@@
Our contributions. We start by devising several algorithms for the GR problem with a CC oracle. A summary of the upper bounds is shown in Table 1. We present results parameterized by four graph parameters: number of edges $m$, maximum degree $\Delta$, treewidth $k$, and degeneracy $d$. With all these parameters, we provide adaptive algorithms (both deterministic and randomized) whose query complexity is polynomial with respect to the parameter and (poly)logarithmic with respect to the number of vertices. When the maximum degree of the hidden graph is known, we also provide a non-adaptive algorithm. The algorithms for $m$, $\Delta$, and $k$ can be combined, giving a randomized adaptive algorithm that uses at most $\mathcal{O}(\min\{m/\log m,\Delta^{2},k^{2}\}\cdot\log n)$ queries (Theorem 3.27). Our algorithm for graphs with bounded degeneracy uses $\mathcal{O}(d^{2}\log^{2}n)$ queries, so it is mildly worse in terms of $n$. However, the degeneracy of a graph is always the smallest one of all the considered parameters, and so this algorithm is applicable to many graph families for which the other algorithms require polynomially many queries in $n$. For example, planar graphs have degeneracy at most five [20] but can have unbounded maximum degree and treewidth.
We also study lower bounds; a summary is shown in Table 2. Black et al. [9] showed that it is not possible to design efficient non-adaptive algorithms for the #\#$C$ queries, and we note that their result applies also for CC queries. In other words, non-adaptive algorithms require $\Omega(n^{2})$ CC queries in the worst case if no additional information is given (Corollary 4.3), even if we know that the graph has treewidth $2$ and thereby has at most $\mathcal{O}(n)$ edges. There is a peculiar observation that very little adaptivity can help. Algorithm 4 has two rounds of adaptivity and query complexity $\mathcal{O}(k^{2}\log n)$ implying that adding one round of adaptivity reduces the dependence on $n$ from $n^{2}$ to $\log n$. Furthermore, we show that no adaptive algorithm can solve the problem for all instance by performing $o(m)$, $o(\Delta^{2})$, $o(k^{2})$, or $o(d^{2})$ CC queries (Theorem 4.1). Thus, our upper bounds are within a (poly)logarithmic factor in $n$ from the lower bound.
Table 1: Summary of our upper bounds. The first corollary follows directly from Black et al. [9].
@@
Theorem 3.33 $\mathcal{O}(d^{2}\log^{2}n)$ yes yes -
Theorem 3.33 $\mathcal{O}\big{(}d^{3}\log n\log(n/d)\big{)}$ no yes existence result
Table 2: Summary of our lower bounds, where the Complexity column states that there is an instance on which $\Omega(\cdot)$ CC queries are needed. A lower bound for randomized algorithms holds for deterministic algorithms, and a lower bound for adaptive algorithms holds for non-adaptive algorithms. The proof of Corollary 4.3 is identical to the lower bound proof for the #\#$C$ oracle [9].
@@
Theorem 4.1 $\Omega(d^{2})$ yes yes -
Corollary 4.3 $\Omega(n^{2})$ yes no even if $k=2$ and $m=\mathcal{O}(n)$
Finally, we compare the CC oracle with other oracles. First, we show that getting explicit information about the connected components rather than just their number [9] leads to more efficient algorithms (Theorem 5.1). Next, we consider the maximal independent set (MIS) oracle introduced by Konrad et al. [18]. While there are instances where the MIS oracle needs $\mathcal{O}(n)$ queries but $\mathcal{O}(\log n)$ queries are sufficient for the CC oracle (Theorem 5.3), these two oracles seem to be roughly equally powerful in the sense that both upper and lower bounds for $m$ and $\Delta$ are nearly the same for both types of oracles. However, CC oracles and MIS oracles are different and neither of them can be efficiently simulated with the other (Theorems 5.5 and 5.7). We also compare the CC oracle and the separation oracle, showing that simulating the former oracle with the latter one requires a quadratic number of queries in the size of the queried set (Theorem 5.11).
@@ 2 Notation and Preliminaries
The subgraph of $G$ induced by $S\subseteq V$, denoted by $G[S]$, has the vertex set $S$ and the edge set $\big{\{}uv\in E(G)\colon u,v\in S\big{\}}$. A subset of vertices is an independent set if the subgraph induced by it has no edges. An independent set is maximal if none of its proper supersets are independent. The degeneracy of a graph $d=d(G)$ is the smallest integer such that every induced subgraph has at least one vertex with degree at most $d$.
@@ 3.1 Bounded Number of Edges
Recently, Black et al. [9] showed an algorithm utilizing an oracle that returns the number of connected components for the subgraph induced by the query set $S$ can reconstruct any hidden graph on $m$ edges with $\mathcal{O}\big{(}m\log(n)/\log(m)\big{)}$ queries. Since our oracle can trivially simulate theirs, the same upper bound follows.
@@
There is an adaptive deterministic algorithm that solves GR with $\mathcal{O}\left(\frac{m\log(n)}{\log(m)}\right)$ CC queries.
@@ 3.2.1 Derandomization
An $(n,\kappa,\kappa^{2})$-splitter of size $\kappa^{\mathcal{O}(1)}\log n$ can be constructed in time $\kappa^{\mathcal{O}(1)}n\log n$.
@@ 3.3 Bounded Treewidth
Algorithm 2 performs the first step: It outputs a supergraph of the hidden graph. Note that, to get a correct result, one does not actually need to know the treewidth of the hidden graph $G$ but an upper bound of it. However, the number of queries conducted increases if the bound is not tight. Next, we will prove the correctness and complexity of Algorithm 2.
@@ 5   Remove edges uvuv for which $v$ and $u$ are in different components from $G^{+}$;
We take advantage of the following fact about existence of small balanced separators in bounded treewidth graphs (see, e.g., Lemma 7.19 of the textbook of Cygan et al. [12]):
@@
A coloring of a graph $G$ is a mapping from the vertex set to some number of colors such that neighboring vertices have always different color. Graphs with small treewidth can be colored with few colors (see, e.g., Sopena [29]):
@@ 3.3.2 Unknown $k$
There exists an adaptive randomized algorithm that solves GR with $\mathcal{O}(\min\{m/\log m,\Delta^{2},k^{2}\}\cdot\log n)$ CC queries. The output of the algorithm is correct with high probability.
@@
We use the adaptive randomized algorithms from Corollary 3.1, Theorem 3.15, and Theorem 3.25. If ran separately, they would stop after performing some number of queries—say, $t_{1}$, $t_{2}$, and $t_{3}$, respectively. Consider instead running the algorithms in parallel, such that in the $i$th iteration we perform the $i$th query of each of the three algorithms. If any of them would stop after performing that query, we output the graph computed by that algorithm. In total, we thus perform $3\cdot\min\{t_{1},t_{2},t_{3}\}=\mathcal{O}(\min\{m/\log m,\Delta^{2},k^{2}\}\cdot\log n)$ CC queries. Further, since the output of each algorithm is correct with high probability, so is the output of this algorithm.
3.4 Bounded Degeneracy
Next, we prove a polylogarithmic upper bound for the query complexity for graphs of bounded degeneracy. The bound holds also for graphs with bounded number of edges, maximum degree, and treewidth, since such graphs also have bounded degeneracy. For planar graphs, the degeneracy $d$ is at most five [20] but the number of edges, the maximum degree and the treewidth can be unbounded in terms of $d$, so the polylogarithmic upper bound holds for a strictly larger family of graphs than with the previous algorithms. However, this comes at the expense of an additional $\mathcal{O}(\log n)$ factor in the query complexity.
Note that if the degeneracy is $d$, then the average degree of every induced subgraph is at most $2d$. Consequently, if the average degree of a graph is at most $2d$, then at least half of its vertices have degree at most $4d$. Our approach is then to repeatedly remove all vertices of degree at most $4d$ from $G$, at least halving the number of vertices in each iteration.
Theorem 3.29.
Algorithm 5 is an adaptive randomized algorithm that solves GR with $\mathcal{O}(d^{2}\log^{2}n)$ CC queries when $d$ is known. The output of the algorithm is correct with high probability.
Proof 3.30.
Analogously to the proof of Theorem 3.2, one can show that after running Algorithm 1 with a possibly incorrect value $4d$ for the maximum degree, the outputted graph $G^{\prime}$ has an edge $uv$ if and only if the hidden graph has that edge with high probability for all vertices $v$ with degree at most $4d$ in $G$. In other words, with high probability, we correctly identify all neighbors of vertices with degree at most $4d$ in $G$. Denote this set of vertices by $L$. It then remains to identify the structure of the subgraph of the hidden graph induced by $V\setminus L$. This is precisely what Algorithm 5 does, repeating the process until we have identified the neighbors for all vertices. Since at least half of the vertices of each induced subgraph have degree at most $4d$, this process gets repeated at most $1+\log_{2}n$ times. Taking into account the query complexity of Algorithm 1, we perform $\mathcal{O}(d^{2}\log^{2}n)$ CC queries in total. Finally, recall from the proof of Theorem 3.2 that Algorithm 1 works correctly with probability at least 11/n1-1/n. Since $(1-1/n)^{1+\log_{2}n}$ tends to $1$ as $n$ increases, Algorithm 5 works correctly with high probability.
Input: A vertex set $V$, a CC oracle $\operatorname{CC}(\cdot)$, degeneracy $d$
Output: The hidden graph $G$ with high probability
1 Initialize $G^{+}$ to a complete graph on $V$;
2 $V^{\prime}\leftarrow V$;
3 while $V^{\prime}\neq\emptyset$ do
4   Let $G^{\prime}$ be the output of Algorithm 1 on the vertex set $V^{\prime}$ with (incorrect) maximum degree $4d$;
5   Remove edges $uv$ from $G^{+}$ for which edge uvuv is not in $G^{\prime}$ and $u,v\in V^{\prime}$;
6   Let $L$ be the set of vertices in $G^{\prime}$ with degree at most $4d$;
7 $V^{\prime}\leftarrow V^{\prime}\setminus L$;
8
9return $G^{+}$;
Algorithm 5 An algorithm for GR for graphs with bounded degeneracy
3.4.1 Derandomization
If we instead apply the non-adaptive deterministic algorithm from Theorem 3.13 instead of the randomized algorithm for bounded degree graphs, we get an adaptive deterministic algorithm for graphs with bounded degeneracy.
Theorem 3.31.
There exists an adaptive deterministic algorithm that solves GR with $\mathcal{O}\big{(}d^{3}\log(n)\log(n/d)\big{)}$ CC queries when $d$ is known.
Proof 3.32.
The proof is analogous to the proof of Theorem 3.29 except that we use a $d$-Query-Scheme similarly to Theorem 3.13 to eliminate randomness.
3.4.2 Unknown $d$
The idea for adjusting the algorithm for graphs of unknown parameter value works similarly to before, but some care is needed to identify that our guess for the degeneracy was too low.
Theorem 3.33.
There exist an adaptive randomized algorithm and an adaptive deterministic algorithm that solve GR with $\mathcal{O}\big{(}d^{2}\log^{2}n\big{)}$ and $\mathcal{O}\big{(}d^{3}\log(n)\log(n/d)\big{)}$ CC queries, respectively. The output of the randomized algorithm is correct with high probability.
Proof 3.34.
We start by guessing that $D=1$ is an upper bound for $d$, and then run Algorithm 5 or the algorithm from Theorem 3.31 as if the known value of $d$ were $D$. If we use the former algorithm and it runs correctly or we use the latter algorithm, then we identify that at least half of the vertices have degree at most $4(D+1)$ in each iteration if $D\geq d$. This may also occur if $D<d$, but if does not occur, then we know that $D<d$ and we double the value of $D$ and restart the algorithm. The query complexity is again dominated by the last round, for which $D<2d$.
It remains to show the high probability of the correctness of the randomized algorithm. Recall that the probability of Algorithm 5 succeeds is at least $(1-1/n)^{1+\log_{2}n}$, and note that if the algorithm succeeds on all $\mathcal{O}(\log d)$ rounds of adaptivity, then the output is correct. The probability of this is at least $(1-1/n)^{\mathcal{O}(\log^{2}n)}\approx(1/\mathrm{e})^{\mathcal{O}(\log^{2}(n)/n)}$, which tends to $1$ as $n$ increases.
@@ 4 Lower Bounds
In this section, we show that the query complexity of our randomized algorithms cannot be improved by a factor larger than $\mathcal{O}(\log n)$ with a proof similar to the lower bound proof of Konrad et al. [18]. Whilst Black et al. [9] have showed a tighter upper bound with respect to $m$ for the #\#$C$ oracle, their information-theoretical argument does not work here because our oracle gives more bits of information than theirs. However, we confirm that their non-adaptive lower bound is applicable for the CC oracle, showing that non-adaptive algorithms require $\Omega(n^{2})$ CC queries if they are not given additional information about the structure.
@@
No algorithm can solve GR with $o(m)$, $o(k^{2})$, $o(\Delta^{2})$, or $o(d^{2})$ CC queries with probability greater than $1/2$.
@@
Suppose now that there is a randomized algorithm that outputs the correct graph for each hidden graph $G\in\mathcal{G}_{\eta}$ with probability greater than $1/2$ while performing at most $\eta^{2}-1$ queries. By Yao’s lemma, there is then a deterministic algorithm that performs $\eta^{2}-1$ queries and outputs the correct graph for more than half of the hidden graphs $G\in\mathcal{G}_{\eta}$. However, observe that for any CC query on a set $S\subseteq V$, there are exactly two possible outputs over the graphs $G\in\mathcal{G}_{\eta}$: the vertices in $S\setminus(C_{1}\cup C_{2})$ are always in connected components of size $1$ and the remaining connected components induced by $S$ are either $S\cap C_{1}$ and $S\cap C_{2}$ or just $S$. Consequently, the deterministic algorithm can identify in $\eta^{2}-1$ queries only $2^{\eta^{2}-1}\leq|\mathcal{G}_{\eta}|/2$ hidden graphs, resulting in a contradiction. Thus, $\Omega(\eta^{2})$ queries are needed. We have $k,\Delta,d\leq 2\eta$ and $m\leq 2\eta^{2}$ for this family of hidden graphs, completing the proof.
Corollary 4.3 ([9]).
@@
5 Relationship with Other Oracles
In this section, we compare our CC oracle to the connected component counting oracle, maximal independent set oracle, and the separation oracle.
5.1 Relationship with the Connected Component Counting Oracle
We start by proving that it is not sufficient to use the connected component counting (#\#$C$) oracle of Black et al. [9] for obtaining our results, since $\Omega(n/\log n)$ #\#$C$ queries are needed for already very constrained graph families. The main observation for this result is that the #\#$C$ oracle has $n+1$ possible distinct values it can output, and so in $t$ queries we can correctly identify only $(n+1)^{t}$ graphs. By constructing a family of graphs larger than that, we can show that more than $t$ queries are needed. Similar combinatorial arguments were used by Michel and Scott [22] and Black et al. [9] to prove lower bounds for the maximal independent set oracle in bounded degree graphs and for the #\#$C$ oracle in graphs with $m$ edges.
Theorem 5.1.
No algorithm can solve all instances of GR with maximum degree, treewidth, or degeneracy at most $p$ with $\Omega(np/\log n)$ #\#$C$ queries with probability greater than $1/2$.
Proof 5.2.
Without loss of generality, assume that $p$ divides $n$. Let $G$ be a graph on $n$ vertices $v_{1},v_{2},\dots,v_{n}$ where there is an edge $v_{i}v_{j}$ if and only if $\lfloor(i-1)/p\rfloor=\lfloor(j-1)/p\rfloor$ or i=j1i=j-1. The maximum degree, treewidth, and degeneracy of $G$ are all at most $p$.
We can now construct a family of graphs $\mathcal{G}$ by considering all subgraphs of $G$ that retain all edges of the form $v_{i}v_{i+1}$. The size of the family is then
$|\mathcal{G}|=2^{\binom{p-1}{2}\cdot\frac{n}{p}}=(n+1)^{\binom{p-1}{2}\cdot\frac{n}{p}\cdot\log_{n+1}(2)}.$
Suppose now that there is a randomized algorithm that outputs the correct graph for each hidden graph $G^{\prime}\in\mathcal{G}$ with probability greater than $1/2$ while performing at most $t\coloneqq\binom{p-1}{2}\cdot\frac{n}{p}\cdot\log_{n+1}(2)-1$ #\#$C$ queries. Then, by Yao’s lemma, there is a deterministic algorithm that performs $t$ and outputs the correct graph for more than half of the hidden graphs $G^{\prime}\in\mathcal{G}$. However, a deterministic algorithm can only have $(n+1)^{t}\leq|\mathcal{G}|/(n+1)$ distinct outputs if it performs $t$ queries, and so the output is incorrect for more than half of the hidden graphs.
5.2 Relationship with the Maximal Independent Set Oracle
We next compare the CC oracle against the maximal independent set (MIS) oracle of Konrad et al. [18], which returns a maximal independent set of $G[S]$. Note that the MIS oracle can be adversarial, that is, it may return the “least useful” maximal independent set. We start this section by demonstrating that CC queries are more powerful than MIS queries on some instances.